the vapour density of pclis 104.16 but when heated to 230.c its vapour density is reduced to 62. the degree os dissociation of pcl5 at this temperature will be
 

Grade:12th pass

1 Answers

Arun
6 years ago
as you know , i = 1+ (y-1)α...α = degree of dissociation
y= no of moles of products from one mole reactant
for PCl5 , y = 2
=> i = 1+α
also, y = (normal v.d) /(abnormal v.d.) = 104.16 / 62 = 1.68
= >α = 0.68

Regards
Arun (askIITians forum expert)